We expect to sell 200 units per year at $1 net cash flow each into perpetuity (i.e., we expect net cash flows of $200 per year in perpetuity). In one year, we will know more about the project. If it looks successful, sales projections will be revised upward to 250 units per year and, if the project looks like a failure, sales projections will be revised downward to 150 units per year. Assume that success and failure are equally likely. The project costs $900 and the discount rate is 20%. If we decide to abandon the project at the end of the first year, we can sell it for $800.

Use the decision tree method to determine the NPV of the project.
Answer #1

In the present scenaario we have to cases. In the first case, the project is continued indefinitely whereas in the second case the project can be abandoned. It is discussed in the following ways:

CASE 1: Net Present Values = Cash Inflows - Cash Outflows = 200/.2 (Cash flows for perpetuity) - 900 = $100

Cash flows for perpetuity is calculates as an equal weight of percentage is given to both, if it successful or unsuccessful. The cash flows when the project is successful, there is $250*0.5=$125 and then the cash flows when the project when the project is not successful, there is $150*00.5=$75. The summation would be $200.

CASE 2: Net Present Values = Cash Inflows - Cash Outflows = 200/1.2( for the first year) +800/1.2 - 900 = -$66.67.

Hence, it is advisable that the company does mot disband the project afte one year